OpenStudy (anonymous):

evaluate the integral dx/sqroot x(1+x)

OpenStudy (amistre64):

\[\int\frac{dx}{\sqrt{x(1+x)}}\] ?

OpenStudy (anonymous):

si :)

OpenStudy (anonymous):

the sq root is only for x

OpenStudy (amistre64):

decompose it maybe since its already factored?

OpenStudy (amistre64):

\[\frac{1}{\sqrt{x}(1+x)} = \frac{A}{\sqrt{x}}+\frac{B}{(1+x)}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what u do?

OpenStudy (amistre64):

you solve for A and B by reworking the problem back into the left side; \[1 = A(1+x) + B(\sqrt{x})\] when x=-1; 1 = 0 + Bsqrt(x); B = 1/sqrt{x} when x = 0; 1 = A(1+0) + 0 ; A = 1

OpenStudy (amistre64):

that dint work ... missed the sqrt{-1} :)

OpenStudy (amistre64):

my integral tables says: \[\int\frac{dx}{(1+x)\sqrt{x}}=2\ \tan^{-1}\left(\sqrt{x}\right)\]maybe?

OpenStudy (amistre64):

tan-1(sqrt(x)) derives to what? \[\frac{1}{1+\sqrt{x}^2}*\frac{2}{2\sqrt{x}}\] it works :)

OpenStudy (anonymous):

i did not understand

OpenStudy (amistre64):

since tan^-1(x) derives down to get 1/(1+x^2) ; then integrals with the same structure will suit back up ....

OpenStudy (amistre64):

just as x^3 derives down to 3x^2 3x^2 integrates back up to x^3 +c

OpenStudy (amistre64):

the derivate and the integral are inverse functions

OpenStudy (anonymous):

I think you make a substitution of u=x^(1/2) and go from there.

OpenStudy (amistre64):

aint seen loki in awhile :)

OpenStudy (anonymous):

And you've taken over the world in the meantime.

OpenStudy (amistre64):

it was either that or .... well there wasnt another option lol

OpenStudy (anonymous):